Sparks Company reported sales revenue of $360,000, operating expenses of $50,000, and a net loss of $45,000 for the most recent fiscal year. What amount did Sparks report for cost of goods sold?
Answer #1

Let cost of goods sold be x.

Gross profit/loss=Sales Revenue-COGS

=3,60,000-x .......................................(i)

Net Loss= Gross Profit- operating Expenses

   -45,000 =(3,60,000-x)-50,000 (From (i) above)

x=3,60,000-50,000+45,000

x=3,55,000

So,COGS is $ 3,55,000.